OpenStudy (anonymous):

Find the limit of the trigonometric function. Limit as x approaches 0 is sinx(1-cosx)/x

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{ sinx }{ x }*(1-cosx)= 1*(1-1)=0\]

OpenStudy (anonymous):

why does (1-cosx) become 1

OpenStudy (anonymous):

it doesn't , it becomes (1-1)

OpenStudy (anonymous):

ok but then why does cosx become 1-1

OpenStudy (anonymous):

okay do you know that \[\lim_{x \rightarrow 0}\frac{ sinx }{ x }=1\] or not??

OpenStudy (anonymous):

yes i do get that

OpenStudy (anonymous):

because cos0 =1 and 1-1=0

OpenStudy (anonymous):

ohhhh ok then..... but lets say the bottom of the x under the sinx was squared would it still give me the same answer

OpenStudy (anonymous):

let's see about that .. \[\frac{ sinx }{ x }*\frac{ (1-cosx)(1+cosx) }{x(1+cosx) }\]

OpenStudy (anonymous):

\[\frac{ sinx }{ x }\frac{ \sin^2(x) }{ x(cosx+1) }=\frac{ \sin^2(x) }{ x^2 }*\frac{ sinx }{ cosx+1 }\]

OpenStudy (anonymous):

\[=1^2*\frac{ 0 }{ 2 }=0\]

OpenStudy (anonymous):

since both of them gave me zero can it be done by just placing the \[x^{2}\] underneath the \[\sin ^{2}x\] and solving it the same way you did the first time

OpenStudy (anonymous):

yes , actually \[\lim_{x \rightarrow 0} \frac{ 1-cosx }{ x }=0\] is a rule , but I thought it's better to show you how it comes

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

welcome